If you had $8,400 placed in an account that earned 5 ½ percent interest, paid just once a year, how much money would be in the account at the end of four years, assuming you made no withdrawals of any kind from the account?
Answer by godeapurva(7)

Principal= $8,400
Rate of interest = 5 ½ % = 5.5%
Years = 4
Simple interest = P*N*R/100
Where P= Principal, N=Years & R= Rate of interest
Thus 8400*5.5*4/100
= 1848
Thus the amount to be present in the account after 4 years = P+ SI
= 8400 + 1848
= 10248 $